# Towards a universal QAOA protocol: Evidence of a scaling advantage in solving some combinatorial optimization problems

### v2

@@ III.3 Experiments
In this section, we show numerically and experimentally how noise affects LR-QAOA. Before moving to numerical simulations of LR-QAOA under depolarizing noise, we want to show LR-QAOA’s ability to correct errors along the way. In Fig  6-(a), the noiseless evolution of the eigenvalues of the cost Hamiltonian for LR-QAOA is presented. In Fig.6-(b) the same protocol is shown but this time depicts the evolution under full inversion of the qubits using a layer of X gates applied at $p=15$. At $p=16$, the eigenvalues experience a full inversion of probabilities with high energy bitstrings now having a large probability. This is quickly corrected by LR-QAOA increasing the probability of getting the optimal solution. This inversion comes with the price of a reduction in the success probability from $96.1\%$ in (a) to $28.5\%$ in (b).
@@
Figure 7 shows 10-qubit W-Maxcut problem results using LR-QAOA and depolarizing noise for $\lambda$s of Eq. 14 from $10^{-1}$ to $10^{-5}$ affecting only the 2-qubit gates. We compare these results with the ones obtained using ionq_aria (diamonds) for the same problem. Fig. 7-(a) shows how the probability of success vs. the number of LR-QAOA layers for the different depolarizing noise strengths. Fig. 7-(b) shows the approximation ratio vs. the number of LR-QAOA layers for the same problem. In both plots, ionq_aria results agree with a $\lambda=6\times 10^{-4}$. Finally, Fig. 7-(c) shows the maximum probability of success vs. $\lambda$. At $\lambda>5\times 10^{-2}$ information can still be recovered and at $\lambda=10^{-5}$ errors are so slowly added that the algorithm can in some sense correct them, similar to what we show in Fig. 6-(b).
Figure 7: Results of LR-QAOA for the 10-qubit random W-MaxCut problem at different $p$ when noise is added at different depolarizing noise strength $\lambda$. (a) probability of success (b) approximation ratio (c) largest probability of success for each $\lambda$. Diamonds represent the same problem executed on ionq_aria. Different lines represent the strength $\lambda$ of the depolarizing channel. The shaded line shows the probability of obtaining the optimal solution by random guessing.
Figure  8-(a) shows the probability of success vs. the number of LR-QAOA layers of random cases of the W-MaxCut for variables from 5 up to 30 running on ibm_brisbane. We use 10000 samples for each problem size. We do not include information for larger problem sizes because no optimal solution is observed for them. The dashed line represents the probability of success of mitigated samples of a random sampler. In other words, circles above the dashed line of its respective color cannot be explained as the result of a random process and therefore can be attributed to LR-QAOA. To contextualize our outcomes, observing the optimal solution for the 30 qubits problem with a random sampler requires in the worst case $536.870.9122^{30}/2=$ 536.870.912 evaluations of the cost function. In our experiment, we find the optimal solution 2 times at 13 layers using LR-QAOA on a noisy device using $10000$ samples and the mitigation technique. This means $10000\times 30$ further evaluations representing an improvement over random guessing of $536.870.912/310.000\approx 1732$ times.
Figure  8-(b) shows the approximation ratio of the instances of W-MaxCut from 5 to 109 qubits using LR-QAOA on ibm_brisbane. The vertical dashed line at $p=10$ indicates the number of layers for which the best performance of LR-QAOA is obtained. After $p=10$, the system is slowly moved towards a maximally mixed state. At $p=100$, it is reached in all the cases. We attribute this phenomenon to the characteristics of LR-QAOA to correct errors on the way. This leads to an interesting behavior, for instance, at $p=3$ the approximation ratio is the same as that at $p=40$ for the 109-qubit case, despite the latter requiring roughly 13 times more time and 2-qubit gates than the former.
Figure 8-(c) shows the maximum probability over all layers vs. the number of qubits for the 1D W-MaxCut experiment. The dashed line that represents the quadratic speedup $1/2^{N_{q}/2}$ is added as a reference. In the experiments, the highest probability occurs within the range of $p=10$ to $13$. The experiments hold a similar decay to the quadratic speedup, with a shift that can be attributed to the mitigation technique. Additionally, Fig. 8-(d) shows the best approximation ratio among all the samples vs. the number of qubits. The maximum average approximation ratio for the 109-qubit experiment is $r=0.64$ with the best sample having a $r=0.84$.
@@
Figure 8: Experimental results of the LR-QAOA protocol on ibm_brisbane for W-MaxCut problems. (a) Probability of success vs. number of layers of the LR-QAOA protocol. Colors represent the number of qubits from 5 to 30 qubits. The dashed lines ’random limit’ represent the success probability of a random sampler after the mitigation technique is applied for each problem with the same color. (b) Approximation ratio vs. LR-QAOA layers. Colors represent problems from 5 to 109 qubits. (c) Best probability of success of (a) vs. number of qubits. The dashed line representing a quadratic speedup is added as a reference. (d) Best approximation ratio of all the samples in the experiment vs. number of qubits.
Figure  9-(a) shows the approximation ratio of the 109 qubits W-MaxCut problem using LR-QAOA from $p=3$ to $p=100$. At $p=10$, the maximum approximation ratio is reached for the three devices ibm_brisbane, ibm_kyoto, and ibm_osaka. The noise at larger $p$ leads the system towards a maximally mixed state, so we include the dashed line that represents the approximation ratio $r=0.5326$ of a random sampler after the mitigation technique is applied. Unexpectedly, at $p=100$, results for ibm_kyoto and ibm_osaka still deviate from the random sampler and therefore some information of the LR-QAOA protocol is present. The circuit used requires 21200 CNOT gates and a total time of $\approx 132\mu s$. This is an indication of the resilience of the LR-QAOA to noise.
At this scale, we surpass the point where exact classical simulation of LR-QAOA is feasible both in terms of the number of qubits and depth of the circuit. Therefore, we believe that this experiment can be presented as a quantum supremacy/utility experiment. This means that a classical algorithm cannot mimic the sampling properties of LR-QAOA at large $p$ and $N_{q}$. There are different techniques proposed for addressing the simulation of quantum supremacy [48] or utility [49] experiments after their publication (e.g., [50, 51]), so the validation of this remains subject to evaluation within the research community. Independently of the answer, these results indirectly imply the efficacy of LR-QAOA to solve COPs in scenarios involving more than 42 qubits.
@@
Figure 9: (a) 109-qubit W-MaxCut experiment on IBM Eagle devices using LR-QAOA. The upper-left plot represents the best approximation ratio observed from the 10000 samples at each $p$. The bottom-left plot shows the average approximation ratio where the dotted line represents the raw result from ibm_brisbane and the solid lines represent the mitigated results over the three different IBM devices. The black dashed line is the limit where the system reaches the maximally mixed state. The right plot shows the IBM Eagle layout with the 1D random W-MaxCut on it. The colors represent the random weights chosen from 4 possible values 0.1, 0.2, 0.5, and 2.0 for each edge in the graph. The inset shows how the 2-qubit gates are implemented using only a depth of 2 for each LR-QAOA layer. (b) 25-qubit W-MaxCut experiment comparative analysis for 3 different vendors. The bottom-left plot presents the graph of the W-MaxCut selected with the corresponding weight values. In the upper plot, the approximation ratio of the devices with the triangles, circles, diamonds, and pentagons correspond to the mitigated results from a noiseless simulation, ibm_brisbane, ionq_aria, and quantinuum_H2, respectively. Because of a limitation in the maximum number of single-qubit and two-qubit gates, there are no results above $p=15$ using ionq_aria. In the bottom right plot the probability of success for the same experiment.
@@
Figure  9-(b) presents a comparative analysis between ionq_aria, quantinuum_H2, and ibm_brisbane in solving a 25-qubit instance of the W-MaxCut problem. The number of samples is 10000 for the noiseless simulator and ibm_brisbane, 1000 for ionq_aria, and 50 for quantinuum_H2. The performance of quantinuum_H2 stands out, achieving a maximum approximation ratio of $r=0.95$ at $p=50$, compared to $r=0.98$ at $p=50$ of the noiseless simulator, ibm_brisbane’s $r=0.80$ at $p=10$, and ionq_aria $r=0.90$ at $p=10$.
@@
From a time perspective, executing an instance of W-MaxCut LR-QAOA for $p=10$ on ibm_brisbane requires approximately $\approx 13.2\mu s$, whereas ionq_aria completes the same task in about $\approx 144ms$, and quantinuum_H2 in $\approx 36ms$. The three devices successfully identify the optimal solution for this problem, with quantinuum_H2 achieving a maximum probability of success of 0.10 at $p=75$, ionq_aria achieving a maximum probability of success of 0.008 at $p=10$, and ibm_brisbane reaching 0.0005 at $p=10$. This means that quantinuum_H2 is 12.5 times more effective in finding the optimal solution than ionq_aria, and 200 than ibm_brisbane.
@@ IV Conclusions
In this work, we have presented numerical and experimental evidence that LR-QAOA constitutes a universal QAOA protocol. This means that this protocol works efficiently independently of the problem tested, increasing the probability of success as the number of layers increases. We simulate MIS, BPP, TSP, MaxCut, W-MaxCut, 3-MaxCut, KP, PO, Max-2-SAT, and Max-3-SAT problems with up to 42 qubits and 400 layers on the modular supercomputer JUWELS. Additionally, we test LR-QAOA using W-MaxCut problems from 5 to 109-qubit cases and $p$ from 3 to 100 on real quantum hardware using ibm_brisbane, ibm_osaka, ibm_kyoto, quantinuum_H2, and ionq_aria finding that LR-QAOA is resilient to noise. We show that this characteristic comes from the algorithm’s ability to overcome noise. While noise pushes the system through a maximal mixed state, LR-QAOA pushes to a minimum energy correcting errors on the way.
@@
One important conclusion from this work is that one can completely suppress the classical optimization step in QAOA. With the fixed schedule in LR-QAOA, one can inexpensively jump to the number of layers needed to observe the optimal solution within a given number of samples.
@@
We show the evolution of LR-QAOA from the perspective of the amplitudes of the computational basis states. This change in framework allows us to explain the evolution of the amplitudes under the application of $U_{C}$ and $U_{B}$. Under the application of $U_{C}$, each amplitude is rotated proportionally to the state energy. The case of $U_{B}$ is more complex but every amplitude evolves with contributions from the other states’ amplitude with the Hamming distance as the indicator of how to group their contribution. The annealing characteristics of LR-QAOA along with a constant rotation of the states’ amplitude (constant slope of the linear ramp) under $U_{C}$ allow the exploitation of an interference pattern that enhances the optimal solution in the different COPs.
@@
We observe that the success probability of the optimal solutions using LR-QAOA for the different COPs seems to scale as $probability(x^{*})\approx 1/2^{(\eta N_{q}/p)}$ for a constant $\eta$. We add further evidence that this is the case using W-MaxCut, MaxCut, Max-2-SAT, and Max-3-SAT with up to 42 qubits. Using $p=N_{q}$, we find that on average, the $probability(x^{*})$ remains nearly constant for W-MaxCut, MaxCut, and Max-3-SAT. The Max-2-SAT case is an exception, using $p=N_{q}$, it already shows an exponential decay in the probability of success. We think this is a consequence of problems with a high concentration of solutions close to the optimal solution. With all of this information, we conjecture that $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for general problems.
This conjecture can be taken as an estimate of the expected performance of LR-QAOA. It implies that LR-QAOA can solve typical cases of NP-Hard problems in polynomial time. However, we also find particular cases where LR-QAOA does not follow a $probability(x^{*})=\mathrm{cte}$ when $p=N_{q}$. In such cases, there is a sub-exponential speedup. It is still possible that in the general case a $p=\mathrm{Poly}(N_{q})$ holds the $probability(x^{*})=\mathrm{cte}$. However, we suspect that in the limit, i.e., a random oracle with just one optimal solution and the first excited state completely degenerated, a performance similar to that of Grover’s algorithm will be obtained. We also observed that the approximation ratio increases with $p$ and seems constant at a specific $p$ independently of the problem size for MaxCut, W-MaxCut, and 3-MaxCut problems.
We compare SA, LR-QAOA, and CPLEX in terms of TTS for fully connected W-MaxCut problems, observing a scaling improvement in LR-QAOA. We attribute this advantage to the unique characteristics of LR-QAOA. While classical solvers require time to learn the energy structure of the problem, LR-QAOA leverages the $U_{C}$ phenomenon (Eq. 7), which inherently contains the full energy information in the rotation of the amplitudes. This information is then efficiently utilized by the $U_{B}$ gate to amplify the probability of obtaining low-energy solutions.
Moreover, we find that LR-QAOA is resilient to noise. This is important as we are at a stage where quantum computers have moderate noise. We simulate a MIS using LR-QAOA with $p=80$ and at $p=15$ we add a layer of X gates to see how the algorithm evolves under this noise. We find that there is an interplay between noise and the algorithm trying to get to the ground state. If the noise is low enough, the algorithm corrects it to some extent. We use depolarizing noise for the simulation of a 10-qubit W-MaxCut problem using the LR-QAOA protocol and find that it fits well with the experimental results on ionq_aria.
@@
We find that there is an effective number of layers for what the real device shows the best performance. We call it the $p_{\mathrm{eff}}$, this parameter can be used to measure the progress of quantum technology for combinatorial optimization. For IBM Eagle devices and ionq_aria $p_{\mathrm{eff}}=10$ and for quantinuum_H2 is $p_{\mathrm{eff}}=50$ for a 1D topology problem. We expect the $p_{\mathrm{eff}}$ decreases for a fully connected graph problem.
@@
The experimental results make us optimistic that in the near future, we will start to compete against classical solvers to find the best-known solutions for COPs. For example, quantinuum_H2 already shows its peak performance at $p=50$ and loses little performance at $p=100$. At the peak point, the device reaches the best approximation ratio of $r=0.95$ and $probability(x^{*})=0.08$ for a 25-qubit problem. On the other hand, the inaccuracy of ibm_brisbane is still compensated by its sample rate for the same problem. The next step, left for future work, is to test fully connected W-MaxCut problems where quantinuum_H2 and ionq_aria have a clear advantage because they have a fully connected layout.
The 109-qubit experiment already in the regime of quantum supremacy constitutes indirect proof that the LR-QAOA protocol works for larger cases. These cases constitute a set of experiments that cannot be simulated exactly on classical computers, and even approximations will become hard as $p$ grows. We find that some information on the LR-QAOA algorithm remains for the largest case, $N_{q}=109$ and $p=100$.
@@
The datasets for W-MaxCut problems used and/or analyzed during the current study are available from the following publicly accessible repository https://jugit.fz-juelich.de/qip/lr-qaoa. The other datasets are available from the corresponding author upon reasonable request.

### v3

@@ III.3 Experiments
[…]
Figure 6: Overlap between the success probability of LR-QAOA under noise and the ideal probability vs the accumulated error. (a) Noise simulation using depolarizing noise for 8, 10, 12-qubit problems with p=10, 20, 40, 25 depolarizing noise errors between $\lambda=10^{-5}$ and 111, and 3 random graphs with edge density $E_{d}=0.2$, 0.5, and 1. The fitting parameter in Eq. 18 is $k_{0}=1.82$, which corresponds to the black line. (b) Overlap of the success probability on real QPUs, ibm_fez (green) and Quantinuum H1-1E (orange) for 5 to 15 qubits. The errors perceived using the noise model are $\lambda_{H1-1E}=2.5\times 10^{-4}$ and $\lambda_{ibm\_fez}=25\times 10^{-4}$. (c) Overlap probability vs. the number of two-qubit gates for noise models of ibm_fez and H1-1E. The dotted line represents the number of 2-qubit gates where an overlap of 10% is reached.
Figure 6 shows simulations of different LR-QAOA configurations for the WMaxcut varying p, $\lambda$, $N_{q}$, and $E_{d}$ under the depolarizing noise model. We make a distinction in this figure by the number of qubits, but the markers represent cases with different p, $\lambda$, and $E_{d}$ as well. Therefore, even if different parameters could have an impact on the solution, the noise can be well described by the number of 2-qubit gates, the depolarizing noise, and a single fitting parameter. The fitting parameter in Eq. 18 is $k_{0}=1.82$ for WMaxcut; it might be COP dependent, but further analysis is needed.
We use this noisy model on experimental results for ibm_fez and H1-1E. The results are shown in Fig. 6(b) for 5 to 15-qubit problems of fully connected WMaxcut. For the case of ibm_fez, we use the parity twine chains (PTC) [44, 42] strategy to encode the LR-QAOA quantum circuit into a 1D-chain of qubits of the QPU. The number of 2-qubit gates for each layer of LR-QAOA is $N_{2q}=N_{q}(N_{q}-1)/2$ for H1-1E and $N_{2q}=N_{q}^{2}-1$ for ibm_fez. Based on the noise model Eq. 14, the fitted average error rates are $\lambda=2.5\times 10^{-4}$ for H1-1E and $\lambda=25\times 10^{-4}$ for ibm_fez. The $\lambda$ can be interpreted as the average 2-qubit error of the QPU for the LR-QAOA on WMaxcut problems. These errors are similar to the average 2-qubit gate error measured by Randomized benchmarking (RB) [49], which is $36\times 10^{-4}$ for ibm_fez and $9\times 10^{-4}$ for H1-1.
Figure 6(c) shows the estimated overlap probability, $p_{ovl}$, as a function of the number of two-qubit gates, under the noise levels observed in ibm_fez and H1-1E. To achieve an overlap of $p_{ovl}=0.1$ (i.e., 10% of the ideal probability), the maximum number of two-qubit gates should be limited to approximately 733 for ibm_fez and 7331 for H1-1E. This limit is independent of the problem size, and therefore, can be used as an estimate of how many 2-qubit gates one can use until the overlap is too short to observe the optimal solution. Even if the error grows exponentially with the number of 2-qubit gates, the probability of success also grows exponentially as the number of layers grows. Therefore, there is a point where the trade-off between noise and LR-QAOA reaches an equilibrium point, and a maximum probability of success is obtained. A decrement in one order of magnitude in the noise leads to an increment in 1 order of magnitude in the number of 2-qubit gates that can be used, for instance, a $\lambda=2.5\times 10^{-5}$ and expecting an overlap of 10% allows to use $N_{g}=73,310$. In the Supplementary Note 4, we extend the depolarizing noise model study to the IonQ Aria QPU.
@@
Figure 7: Experimental results of the LR-QAOA protocol on ibm_brisbane for WMaxcut problems. (a) Probability of success vs. number of layers of the LR-QAOA protocol. Colors represent the number of qubits from 5 to 30 qubits. The dashed lines ’random limit’ represent the success probability of a random sampler after the mitigation technique is applied for each problem with the same color. (b) Approximation ratio vs. LR-QAOA layers. Colors represent problems from 5 to 109 qubits. The shaded region represents the standard deviation over 10 sets of 1000 shots. (c) Best probability of success of (a) vs. number of qubits. The dashed line representing a quadratic speedup is added as a reference. (d) Best approximation ratio of all the samples in the experiment vs. number of qubits.
Figure  7-(a) shows the probability of success vs. the number of LR-QAOA layers of random cases of the WMaxcut for variables from 5 up to 30 running on ibm_brisbane. We use 10,000 samples for each problem size. We do not include information for larger problem sizes because no optimal solution is observed for them. The dashed line represents the probability of success of mitigated samples of a random sampler (See Sec. II.7). In other words, circles above the dashed line of its respective color cannot be explained as the result of a random process and therefore can be attributed to LR-QAOA. To contextualize our outcomes, observing the optimal solution for the 30 qubits problem with a random sampler requires, in the worst case $2^{30}/2=536,870,912$ evaluations of the cost function. In our experiment, we find the optimal solution 2 times at 13 layers using LR-QAOA on a noisy device using 10,00010,00010 , 000 samples and the mitigation technique (See Sec. II.7 for the mitigation technique). This means 10,000×3010,000$\times$ 3010 , 000 × 30 further evaluations, representing an improvement over random guessing of 536.870.912/310.0001732536.870.912/310.000$\approx$ 1732536.870.912 / 310.000 ≈ 1732 times.
Figure  7-(b) shows the approximation ratio of the instances of WMaxcut from 5 to 109 qubits using LR-QAOA on ibm_brisbane. The vertical dashed line at $p=10$ indicates the number of layers for which the best performance of LR-QAOA is obtained. After $p=10$, the system is slowly moved towards a maximally mixed state. At $p=100$, it is reached in all the cases. We attribute this phenomenon to the nature of LR-QAOA, which initially improves faster than the destructive effects of noise. However, above a particular noise threshold, noise begins to dominate, leading to a monotonic decrease in the quality of the solutions obtained. This leads to an interesting behavior, for instance, at $p=3$ the approximation ratio is the same as that at $p=40$ for the 109-qubit case, despite the latter requiring roughly 13 times more time and 2-qubit gates than the former.
Figure 7-(c) shows the maximum probability over all layers vs. the number of qubits for the 1D WMaxcut experiment. The dashed line that represents the quadratic speedup $2/2^{N_{q}/2}$ over random sampling is added as a reference. In the experiments, the highest probability occurs within the range of $p=10$ to 131313. The experiments hold a similar decay to the quadratic speedup, with a shift that can be attributed to the mitigation technique. Additionally, Fig. 7-(d) shows the best approximation ratio among all the samples vs. the number of qubits. The maximum average approximation ratio for the 109-qubit experiment is $r=0.64$, with the best sample having a $r=0.84$.
@@
Figure 8: (a) 109-qubit WMaxcut experiment on IBM Eagle devices using LR-QAOA. The upper-left plot represents the best approximation ratio observed from the 10000 samples at each $p$. The bottom-left plot shows the average approximation ratio, where the dotted line represents the raw result from ibm_brisbane and the solid lines represent the mitigated results over the three different IBM devices. The black dashed line is the limit where the system reaches the maximally mixed state. The right plot shows the IBM Eagle layout with the 1D random WMaxcut on it. The colors represent the random weights chosen from 4 possible values 0.1, 0.2, 0.5, and 2.0 for each edge in the graph. The inset shows how the 2-qubit gates are implemented using only a depth of 2 for each LR-QAOA layer. The shaded region represents the standard deviation over 10 sets of 1000 shots. (b) 25-qubit WMaxcut experiment comparative analysis for 3 different vendors. The bottom-left plot presents the graph of the WMaxcut selected with the corresponding weight values. In the upper plot, the approximation ratio of the devices with the triangles, circles, diamonds, and pentagons corresponds to the mitigated results from a noiseless simulation, ibm_brisbane, ionq_aria, and quantinuum_H2, respectively. Because of a limitation in the maximum number of single-qubit and two-qubit gates, there are no results above $p=15$ using ionq_aria. In the bottom right plot, the probability of success for the same experiment is shown.
Figure  8-(a) shows the approximation ratio of the 109 qubits WMaxcut problem using LR-QAOA from $p=3$ to $p=100$. At $p=10$, the maximum approximation ratio is reached for the three devices ibm_brisbane, ibm_kyoto, and ibm_osaka. The noise at larger $p$ leads the system towards a maximally mixed state, so we include the dashed line that represents the approximation ratio $r=0.5326$ of a random sampler after the mitigation technique is applied. Unexpectedly, at $p=100$, results for ibm_kyoto and ibm_osaka still deviate from the random sampler and therefore some information of the LR-QAOA protocol is present. The circuit used requires 21200 CNOT gates and a total time of $\approx 132\mu s$. This is an indication of the resilience of the LR-QAOA to noise.
At this scale, we surpass the point where exact classical simulation of LR-QAOA is feasible both in terms of the number of qubits and depth of the circuit. The 1D connectivity of the graph makes the simulation of the LR-QAOA suited for approximation methods based on tensor networks. If the problem connectivity is increased, making it hard to be simulated classically, this experiment might be presented as a quantum utility experiment. This means that a classical algorithm cannot mimic the sampling properties of LR-QAOA at large $p$ and $N_{q}$. There are different techniques proposed for addressing the simulation of quantum supremacy [50] or utility [51] experiments after their publication (e.g., [52, 53]), so the validation of this remains subject to evaluation within the research community. Independent of the answer, these results indirectly imply the efficacy of LR-QAOA to solve COPs in scenarios involving more than 42 qubits.
@@
Figure  8-(b) presents a comparative analysis between ionq_aria, quantinuum_H2, and ibm_brisbane in solving a 25-qubit instance of the WMaxcut problem. The number of samples is 10000 for the noiseless simulator and ibm_brisbane, 1000 for ionq_aria, and 50 for quantinuum_H2. The performance of quantinuum_H2 stands out, achieving a maximum approximation ratio of $r=0.95$ at $p=50$, compared to $r=0.98$ at $p=50$ of the noiseless simulator, ibm_brisbane’s $r=0.80$ at $p=10$, and ionq_aria $r=0.90$ at $p=10$.
@@
From a time perspective, executing an instance of WMaxcut LR-QAOA for $p=10$ on ibm_brisbane requires approximately $\approx 13.2\mu s$, whereas ionq_aria completes the same task in about $\approx 144ms$, and quantinuum_H2 in $\approx 36ms$. The three devices successfully identify the optimal solution for this problem, with quantinuum_H2 achieving a maximum probability of success of 0.10 at $p=75$, ionq_aria achieving a maximum probability of success of 0.008 at $p=10$, and ibm_brisbane reaching 0.0005 at $p=10$. This means that quantinuum_H2 is 12.5 times more effective in finding the optimal solution than ionq_aria, and 200 than ibm_brisbane.
@@ IV Discussion
In this work, we have presented numerical and experimental evidence that LR-QAOA constitutes an effective QAOA protocol. This means that this protocol works efficiently for the problem tested, increasing the probability of success as the number of layers increases. We simulate MIS, BPP, TSP, Maxcut, WMaxcut, 3-Maxcut, KP, PO, Max-2-SAT, and Max-3-SAT problems with up to 42 qubits and 400 layers on the modular supercomputer JUWELS. Additionally, we test LR-QAOA using WMaxcut problems from 5 to 109-qubit cases and $p$ from 3 to 100 on real quantum hardware using ibm_brisbane, ibm_osaka, ibm_kyoto, quantinuum_H2, and ionq_aria finding that LR-QAOA is resilient to noise. We show that this behavior arises from the algorithm’s ability to enhance solution quality at a rate that initially outpaces the accumulation of noise. While the overlap of the probability of success decreases exponentially with the number of 2-qubit gates, it is compensated by an exponential growth in the probability of success up to some p. This explains why the highest probability of success does not show up at the smallest number LR-QAOA layers in the experiments but at some other point, e.g., at p=10 in the ibm_brisbane case.
@@
One important conclusion from this work is that one can completely suppress the classical optimization step in QAOA for some COPs. With the fixed schedule in LR-QAOA, one can reduce the set of parameters to tune to only three $\Delta_{\beta}$, $\Delta_{\gamma}$, and $p$.
@@
We show the evolution of LR-QAOA from the perspective of the amplitudes of the computational basis states. This change in framework allows us to explain the evolution of the amplitudes under the application of $U_{C}$ and $U_{B}$. Under the application of $U_{C}$, each amplitude is rotated proportionally to the state energy. The case of $U_{B}$ is more complex, but every amplitude evolves with contributions from the other states’ amplitudes with the Hamming distance as the indicator of how to group their contribution. The annealing characteristics of LR-QAOA, along with a constant rotation of the states’ amplitude (constant slope of the linear ramp) under $U_{C}$, allow the exploitation of an interference pattern that enhances the optimal solution in the different COPs.
We observe that the success probability of the optimal solutions using LR-QAOA for the different COPs seems to scale as $probability(x^{*})\approx 2^{-\eta(p)N_{q}+C}$ for $\eta(p)$ decreasing with $p$ and a constant $C$. We add further evidence to solve fully connected WMaxcut problems. We create 100 random problems and select the 20 of them that require the highest number of iterations for the SA solver. We compare SA, LR-QAOA, TABU, and CPLEX in terms of TTS for these problems, observing a better scaling in LR-QAOA.
@@
We extend the study to Maxcut, Max-2-SAT, and Max-3-SAT with up to 42 qubits. Using $p=N_{q}$, we find that on average, the $probability(x^{*})$ remains nearly constant for WMaxcut, Maxcut, and Max-3-SAT. The Max-2-SAT case is an exception, using $p=N_{q}$, it shows an exponential decay in the probability of success, still above a quadratic speedup over random guessing. We think this is a consequence of problems with a high concentration of solutions close to the optimal solution.
Moreover, we find that LR-QAOA tolerates noise. This is important as we are at a stage where quantum computers have moderate noise. We simulate a MIS using LR-QAOA with $p=80$, and at $p=15$ we add a layer of X gates to see how the algorithm evolves under this noise. We find that even in this scenario, the error does not expand, and the optimal solution can still be found with high probability. We extend the study of noise using depolarizing noise on an FC WMaxcut with different numbers of qubits, layers, and $E_{d}$, and find that $p_{ovl}$ decreases exponentially with the number of 2-qubit gates. Using the same model, we fit experimental results in a Quantinuum H1-1 emulator and ibm_fez QPU. In both cases, our model fit well with an apparent error of $\lambda=2.5\times 10^{-4}$ for H1-1E and $\lambda=25\times 10^{-4}$ for ibm_fez. These errors allow the execution of 733 gates in the case of ibm_fez and 7331 for H1-1E to have an overlap of 10% with the ideal probability.
@@
We find that there is an effective number of layers for which the real device shows the best performance. We call it the $p_{\mathrm{eff}}$, this parameter can be used to measure the progress of quantum technology for combinatorial optimization. For IBM Eagle devices and ionq_aria $p_{\mathrm{eff}}=10$ and for quantinuum_H2 is $p_{\mathrm{eff}}=50$ for a 1D topology problem. We expect the $p_{\mathrm{eff}}$ decreases for a fully connected graph problem because the number of two-qubit gates per layer grows by $O(N_{q}^{2})$ compared with the 1D case, $O(N_{q})$.
@@
The experimental results make us optimistic that LR-QAOA can keep high performance even in the presence of noise. For example, quantinuum_H2 already shows its peak performance at $p=50$ and loses little performance at $p=100$. At the peak point, the device reaches the best approximation ratio of $r=0.95$ and $probability(x^{*})=0.08$ for a 25-qubit problem. On the other hand, the inaccuracy of ibm_brisbane is still compensated by its sample rate for the same problem.
In a recent study [32], 10 hard COPs for classical solvers were introduced, many of which are sparsely connected; this characteristic can make them suitable to be tackled by LR-QAOA. Between these problems is the MIS, and hard instances show up at sizes with a few hundred qubits. For instance, there is a case with 500 qubits and 6256 edges (See Table 7. R 500 005 1 in [32]) for which the optimal solution is not known. For a p=200 LR-QAOA and based in our perceived scale, we need 1’250,200 2-qubit gates, the noise in this case to reach an overlap of 10% is $\lambda=1.45\times 10^{-6}$ and based on the scaling of Fig. 3 the number of samples needed is around 25,000 to observed the optimal solution. Currently, the noise in the QPUs is 2 orders of magnitude above the level of error needed, and the number of qubits is at most 156.
It might be possible that the noise level required of a fault-tolerant quantum computer (FTQC). A recent effort to estimate the overhead scenario of FTQC has been presented for the 8-SAT problem [54], finding that at some point, QAOA combined with amplitude amplification with the FTQC overhead can still outperform the best classical solver for that problem.
@@
All study data are included in this article and the Supplementary Materials. The datasets for the problems used and/or analyzed during the current study are available from the following publicly accessible repository https://github.com/alejomonbar/LR-QAOA.
Code availability
The code is available from the following publicly accessible repository https://github.com/alejomonbar/LR-QAOA.